wo gliders collide on a frictionless air track that is aligned along the x axis. Glider A has an initial velocity v0 and glider B is initially at rest. After they collide, A has a velocity +12v0 and B has a velocity +32v0. Find the ratio of the gliders' masses, mAmB.
Answer #1

here,

the mass of glider A is mA

the mass of glider B is mB

using conservation of momentum

mA * v0 = mA * vA + mB * vB

mA * v0 = - mA * 12 * v0 + mB * (32 * v0)

13 * mA = 32 * mB

mA /mB = 32/13 = 2.46

the ratio of the gliders' masses is 2.46